# Personalized Image Generation with Large Multimodal Models

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ACM Conference; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
In the era of information overload, individuals are overwhelmed with vast amounts of multimodal content on the Web, underscoring the importance of personalized content delivery (Wu et al., 2023, 2013; Zhu et al., 2024). The predominant approach, personalized content filtering like recommender systems (Wang et al., 2023b; He et al., 2020; Wang et al., 2019), relies on user interaction history and contextual information to infer user preferences and filter existing content. However, this approach is constrained by the limited diversity of available content, rendering it inadequate to fully meet users’ varied content needs (see an example in Figure 1). To address this limitation, generating personalized new content is becoming increasingly important across various domains, including personalized movie posters (Shen et al., 2024), advertisements (Yang et al., 2024b; Vashishtha et al., 2024), music (Dai et al., 2022; Plitsis et al., 2024), and fashion designs (Xu et al., 2024; Yu et al., 2019).
@@
Due to the lack of supervised data, Pigeon adopts a two-stage preference alignment scheme to adapt LMMs to the personalized image generation task. As shown in Figure 2, the first stage assumes that user-interacted history images, despite some noise, still partially reflect implicit user preferences. Given a sequence of user-interacted images, Pigeon treats the last one as the target image and the preceding images as the history images. We then mask the target image as a reference image to construct the user’s multimodal instructions and fine-tune Pigeon to reconstruct the target image based on this user’s history images and multimodal instructions, regulating Pigeon to infer user preference from history. In the second stage, Pigeon generates multiple target images based on the first-stage alignment and ranks them using a preference reward strategy, thus forming pseudo-labeled preference data pairs of “chosen” and “rejected” images. Pigeon is then optimized with the preference data pairs via Direct Preference Optimization (DPO) (Rafailov et al., 2024) to generate more personalized target images, enhancing personalization capabilities.
@@
We validate the effectiveness of Pigeon in two popular scenarios: personalized sticker and movie poster generation. Extensive quantitative evaluation demonstrates that Pigeon outperforms the best baseline in personalization, achieving improvements of 7%~31% while maintaining comparable semantic alignment with the reference image. Notably, human evaluation on Amazon MTurk [^1] https://www.mturk.com/. reveals that, on average, 71% participants rate Pigeon-generated images with superior personalization and semantic alignment. Furthermore, we discuss the versatility of Pigeon extending to more domains such as personalized product images, advertisement, and fashion images in Section 2.3, highlighting Pigeon’s broad applicability and significant economic value. Our code and data are available at https://anonymous.4open.science/r/Pigeon.
@@ 2.1. Task Formulation
Personalized image generation aims to synthesize personalized images tailored to implicit user preferences and explicit multimodal instructions. Formally, given a set of use-interacted history images =$\mathcal{H}=\{\bm{x}_{i}\}_{i=1}^{N}$ and multimodal instructions =$\mathcal{R}=\{\bm{x}_{0},\bm{txt}\}$, where $\bm{x}_{0}$ and $\bm{txt}$ represent the reference image and textual instruction, respectively, this goal is to generate a personalized target image $\bm{x}_{N+1}$ that not only meets user visual preferences but also adheres to multimodal instructions by high-level semantic alignment with the reference image. This task has broad applications in enhancing user experience across various domains, such as generating personalized product images in e-commerce or creating personalized movie posters and video thumbnails on platforms like Netflix and YouTube.
@@ 2.2.1. Mask Generation Module
$\bullet$ Reference-aware history filtering. We create a multi-hot binary mask $\bm{m}_{h}$ to mask the low-score tokens according to the history mask ratio $\alpha_{h}\in[0,1]$. This mask filters out noisy or reference-irrelevant history tokens, yielding the filtered token embeddings for each history image:
(4) $[\bm{\tilde{E}}_{1},\dots,\bm{\tilde{E}}_{N}]=\bm{m}_{h}\odot[\bm{E}_{1},\dots,\bm{E}_{N}],$
where $\bm{\tilde{E}}_{i}$ denotes masked history token embeddings. For gradient backpropagation in this discrete sampling process, the Gumbel-Softmax trick (Maddison et al., 2017) is applied to the non-differentiable binary mask.
@@ 2.2.2. Personalization Module
$\bullet$ Multimodal instructions encoding. When directly utilizing the reference image to guide target image generation, LMMs often duplicate the reference image, failing to effectively incorporate personalized information (see empirical results in Section 3.4.2). This highlights the necessity of extracting high-level semantics from the reference image for image generation. To enrich the semantics of the reference image $\bm{x}_{0}$ and enhance the comprehension of multimodal instructions in LMMs, we utilize a caption model (e.g., BLIP-2 (Li et al., 2023) and LLaVA (Liu et al., 2024a)) to generate a textual description of the reference image, which is then tokenized into textual tokens:
@@ 2.2.5. Inference
2.3. Domain Applications of Pigeon
Pigeon empowers LMMs with the capability to generate personalized images, which is applicable in various scenarios such as personalized stickers on social media platforms like Twitter and personalized movie posters on platforms like Netflix (see demonstration in Section 3.5). Beyond these, we showcase the potential of Pigeon in other representative domains.
@@ 3.1.1. Datasets
We conduct experiments on two publicly available datasets, focusing on sticker and movie poster scenarios: 1) SER30K [^3] https://github.com/nku-shengzheliu/SER30K. is a large-scale dataset of stickers, each categorized by theme and annotated with an associated emotion label; and 2) ML-Latest [^4] https://grouplens.org/datasets/movielens., a benchmark dataset containing user ratings on movies.
For the sticker scenario, we exclude low-quality themes or those with fewer than six stickers, constructing user interaction sequences where each user interacts with a single theme. For the movie scenario, we adopt the small version of the dataset, retaining user interactions with ratings of four or higher, sorted by the timestamps. We apply a sliding window of six interactions, moving one step at a time to create data samples for each user in both scenarios. Each sample treats the first five interactions as the user history images and the last as the target image. We split the samples into training, validation, and testing sets with a ratio of 8:1:1. In the sticker testing set, we randomly select one sticker from a different theme than the user history as the reference image, while in the movie poster scenario, the target image is used as the reference. Dataset statistics are summarized in Table 4 in Appendix.
@@ 3.1.3. Evaluation Metrics
  • Semantic similarity. We adopt CLIP (Radford et al., 2021) and DINO (Oquab et al., 2024), two popular visual feature extraction models, to extract image features from generated and history/reference images, and compute the cosine similarity between them to obtain the CLIP Image Score (CIS) and DINO Image Score (DIS). Additionally, the CLIP Score (CS) measures the similarity between generated images and textual descriptions of the history/reference images. To assess the overall performance, we also calculate a unified F1-score, combining the history CIS and reference CS.
@@ 3.1.4. Implementation Details
All the baselines are tuned with a fixed learning rate of $1e^{-5}$. We implement PMG following its default model designs, while other baselines are implemented with Stable Diffusion XL (Podell et al., 2024) as the image generator for fair comparisons.
In Pigeon, the learning rate is also set to $1e^{-5}$, and the history mask ratio $\alpha_{h}$ is fixed at $0.2$. During inference, we select the optimal reference mask ratio $\alpha_{r}\in\{0.0,0.1,\dots,1.0\}$ for each reference image by averaging the history CIS and reference CS. All experiments are conducted using a single NVIDIA-A100 GPU.
@@ 3.2. Quantitative Evaluation (RQ1)
3.3. Human Evaluation (RQ2)
To assess the qualitative performance of Pigeon in personalization and semantic alignment, we conduct a human evaluation on Amazon MTurk [^5] https://www.mturk.com/., comparing it against Grd and two representative baselines: 1) TI, which exhibits the second-best overall performance in Table 1 , and 2) PMG, designed for personalized image generation. The evaluation adopts binary-choice tests across sticker and movie poster scenarios, each with 50 cases. For personalization, we present five user-interacted history images and the generated images, with the question: “When provided with someone’s five previously liked stickers (movies), please select the next sticker (movie poster) that is more attractive to her/him.” For semantic alignment, we display the reference and generated images with the question: “Which image aligns more closely with the semantics of the reference image?” As shown in Table 2, Pigeon consistently surpasses ($\geq 50\%$) the baselines, even the Grd, in personalization and maintains decent results in semantic alignment with reference images. These findings emphasize its superiority in capturing user preferences from noisy history images and effectively integrating multimodal instructions for image generation, which aligns with the quantitative analysis.
@@ 3.4.3. Effect of two-stage preference alignments
$\bullet$ Stage-2: pairwise preference alignment. We evaluate the effect of the second-stage pairwise preference alignment by comparing the performance after the first and second stages of alignments, as shown in Figure 4(e). Despite a slight decline in semantic alignment, the second-stage preference alignment further enhances personalization. This demonstrates the effectiveness of DPO in aligning the generation process more closely with user preferences, ultimately resulting in more personalized image generation.
@@ 3.5. Case Study
In this section, we present two examples of Pigeon-generated images in sticker and movie poster scenarios, along with four user-interacted history images and one reference image. We compare Pigeon with two competitive baselines, TI and PMG, as shown in Figure 5. In the sticker scenario, Pigeon effectively captures the user’s visual preference for Yoda and integrates it with the high-level semantics of the reference sticker, such as “drinking coffee”, achieving impressive personalization and semantic alignment with the reference image. In the movie poster scenario, Pigeon-generated poster for the movie “Rise of the Planet of the Apes” showcases high semantic alignment with the reference poster by emphasizing an intense central ape figure, evoking a similar sense of power and conflict. Meanwhile, it matches the user’s preference for character-centered movie posters with a dark and dramatic color palette. More examples are provided in Figure 6 and Figure 7 in Appendix.
@@ 4. Related Work
$\bullet$ Personalized Content Filtering. Traditional filtering-based personalized content delivery approaches, such as recommender systems (Wang et al., 2023b; Lin et al., 2024; Bao et al., 2023; Zhao et al., 2024), typically rank existing content based on user interaction history and contextual information, delivering the top-ranked content. However, constrained by the limited diversity of available content, they often fall short of meeting users’ diverse needs (Xu et al., 2024; Yu et al., 2019; Wang et al., 2023a), motivating the emergence of personalized content generation across various domains.
@@ 5. Conclusion and Future Work
Appendix A Appendix
Table 4. Dataset of sticker and movie poster scenarios, where each “sample” consists of user-interacted history images and one reference image.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714843isbn: 979-8-4007-1274-6/25/04ccs: Information systems Recommender systemsccs: Information systems Personalizationccs: Information systems Multimedia content creation
@@ 1. Introduction
In the era of information overload, individuals are overwhelmed with vast amounts of multimodal content on the Web, underscoring the importance of personalized content delivery (Wu et al., 2023, 2013; Zhu et al., 2024). The predominant approach, personalized content filtering like recommender systems (Wang et al., 2023b; He et al., 2020; Wang et al., 2019; Gao et al., 2023), relies on user interaction history and contextual information to infer user preferences and filter existing content. However, this approach is constrained by the limited diversity of available content, rendering it inadequate to fully meet users’ varied content needs (see an example in Figure 1). To address this limitation, generating personalized new content is becoming increasingly important across various domains, including personalized movie posters (Shen et al., 2024), advertisements (Yang et al., 2024b; Vashishtha et al., 2024), music (Dai et al., 2022; Plitsis et al., 2024), and fashion designs (Xu et al., 2024; Yu et al., 2019).
@@
Due to the lack of supervised data, Pigeon adopts a two-stage preference alignment scheme to adapt LMMs for personalized image generation. As shown in Figure 2, the first stage assumes that user-interacted history images, despite some noise, still partially reflect implicit user preferences. Given a sequence of these images, Pigeon treats the last one as the target image and the preceding images as the history images. We then mask the target image as a reference image to construct the user’s multimodal instructions and fine-tune Pigeon to reconstruct the target image based on this user’s history images and multimodal instructions, regulating Pigeon to infer user preference from history. In the second stage, Pigeon generates multiple target images based on the first-stage alignment and ranks them using a preference reward strategy, thus forming pseudo-labeled preference data pairs of “chosen” and “rejected” images. Pigeon is then optimized with the preference data pairs via Direct Preference Optimization (DPO) (Rafailov et al., 2024) to generate more personalized target images, enhancing personalization capabilities.
@@
We validate the effectiveness of Pigeon in two popular scenarios: personalized sticker and movie poster generation. Extensive quantitative evaluation demonstrates that Pigeon outperforms the best baseline in personalization, achieving improvements of 7%~31% while maintaining comparable semantic alignment with the reference image. Notably, human evaluation on Amazon MTurk [^1] https://www.mturk.com/. reveals that, on average, 71% participants rate Pigeon-generated images with superior personalization and semantic alignment. Furthermore, we discuss the versatility of Pigeon extending to more domains such as personalized product images, advertisement, and fashion images in Appendix A, highlighting Pigeon’s broad applicability and significant economic value. Our code and data are available at https://github.com/YiyanXu/Pigeon.
@@ 2.1. Task Formulation
Personalized image generation aims to synthesize personalized images tailored to implicit user preferences and explicit multimodal instructions. Formally, given a set of user-interacted history images =$\mathcal{H}=\{\bm{x}_{i}\}_{i=1}^{N}$ and multimodal instructions =$\mathcal{R}=\{\bm{x}_{0},\bm{txt}\}$, where $\bm{x}_{0}$ and $\bm{txt}$ represent the reference image and textual instruction, respectively, this goal is to generate a personalized target image $\bm{x}_{N+1}$ that not only meets user visual preferences but also adheres to multimodal instructions by high-level semantic alignment with the reference image. This task has broad applications in enhancing user experience across various domains, such as generating personalized product images in e-commerce or creating personalized movie posters and video thumbnails on platforms like Netflix and YouTube.
@@ 2.2.1. Mask Generation Module
$\bullet$ Reference-aware history filtering. We create a multi-hot binary mask $\bm{m}_{h}$ to mask the low-score tokens according to the history mask ratio $\alpha_{h}\in[0,1]$. This mask filters out noisy or reference-irrelevant history tokens, yielding the filtered token embeddings for each history image: $[\bm{\tilde{E}}_{1},\dots,\bm{\tilde{E}}_{N}]=\bm{m}_{h}\odot[\bm{E}_{1},\dots,\bm{E}_{N}],$ where $\bm{\tilde{E}}_{i}$ denotes masked history token embeddings. For gradient backpropagation in this discrete sampling process, the Gumbel-Softmax trick (Maddison et al., 2017) is applied to the non-differentiable binary mask.
@@ 2.2.2. Personalization Module
$\bullet$ Multimodal instructions encoding. When directly utilizing the reference image to guide target image generation, LMMs often duplicate the reference image, failing to effectively incorporate personalized information (see empirical results in Section 3.4.2). This highlights the necessity to extract high-level semantics from the reference image for image generation. To enrich the semantics of the reference image $\bm{x}_{0}$ and enhance the comprehension of multimodal instructions in LMMs, we utilize a caption model (e.g., BLIP-2 (Li et al., 2023) and LLaVA (Liu et al., 2024a)) to generate a textual description of the reference image, which is then tokenized into textual tokens:
@@ 3.1.1. Datasets
We conduct experiments on two datasets, focusing on sticker and movie poster scenarios: 1) SER30K [^3] https://github.com/nku-shengzheliu/SER30K. is a large-scale dataset of stickers, each categorized by theme and annotated with an associated emotion label; and 2) ML-Latest [^4] https://grouplens.org/datasets/movielens., a benchmark dataset containing user ratings on movies. Details regarding data processing and dataset statistics are available in Appendix B.
@@ 3.1.3. Evaluation Metrics
  • Semantic similarity. We adopt CLIP (Radford et al., 2021) and DINO (Oquab et al., 2024) to extract image features from generated and history/reference images, and compute the cosine similarity between them to obtain the CLIP Image Score (CIS) and DINO Image Score (DIS). Additionally, the CLIP Score (CS) measures the similarity between generated images and textual descriptions of the history/reference images. To assess the overall performance, we also calculate a unified F1-score, combining the history CIS and reference CS.
@@ 3.1.4. Implementation Details
All the baselines are tuned with a fixed learning rate of $1e^{-5}$. We implement PMG following its default model designs, while other baselines are implemented with Stable Diffusion XL (Podell et al., 2024) as the image generator for fair comparisons. Detailed hyper-parameter settings and computational overhead of Pigeon are summarized in Appendix C.
@@ 3.2. Quantitative Evaluation (RQ1)
3.3. Human Evaluation (RQ2)
To assess the qualitative performance of Pigeon in personalization and semantic alignment, we conduct a human evaluation on Amazon MTurk [^5] https://www.mturk.com/., comparing it against Grd and two representative baselines: 1) TI, which exhibits the second-best overall performance in Table 1 , and 2) PMG, designed for personalized image generation. The evaluation adopts binary-choice tests across sticker and movie poster scenarios, each with 50 cases. For personalization, we present five user-interacted history images and the generated images, with the question: “When provided with someone’s five previously liked stickers (movies), please select the next sticker (movie poster) that is more attractive to her/him.” For semantic alignment, we display the reference and generated images with the question: “Which image aligns more closely with the semantics of the reference image?” As shown in Table 2, Pigeon consistently surpasses ($\geq 50\%$) the baselines, even the Grd, in personalization and maintains decent results in semantic alignment with reference images. These findings emphasize its superiority in capturing user preferences from noisy history images and effectively integrating multimodal instructions for image generation, which aligns with the quantitative analysis. More detailed information can be found in Appendix D.
@@ 3.4.3. Effect of two-stage preference alignments
$\bullet$ Stage-2: pairwise preference alignment. We evaluate the effect of the second-stage pairwise preference alignment by comparing the performance after the first and second stages of alignments, as shown in Figure 4(e). Despite a slight decline in semantic alignment, the second-stage preference alignment further enhances personalization. This demonstrates the effectiveness of DPO in aligning the generation process more closely with user preferences, ultimately resulting in more personalized image generation. Analysis of different preference reward strategies during the second-stage alignment is presented in Appendix E.
@@ 3.5. Case Study
In this section, we present two examples of Pigeon-generated images in sticker and movie poster scenarios, along with four user-interacted history images and one reference image. We compare Pigeon with two competitive baselines, TI and PMG, as shown in Figure 5. In the sticker scenario, Pigeon effectively captures the user’s visual preference for Yoda and integrates it with the high-level semantics of the reference sticker, such as “drinking coffee”, achieving impressive personalization and semantic alignment with the reference image. In the movie poster scenario, Pigeon-generated poster for the movie “Rise of the Planet of the Apes” showcases high semantic alignment with the reference poster by emphasizing an intense central ape figure, evoking a similar sense of power and conflict. Meanwhile, it matches the user’s preference for character-centered movie posters with a dark and dramatic color palette. More examples are provided in Figure 6 and Figure 7 in Appendix F.
@@ 4. Related Work
$\bullet$ Personalized Content Filtering. Traditional filtering-based personalized content delivery approaches, such as recommender systems (Wang et al., 2023b; Lin et al., 2024; Bao et al., 2023; Zhao et al., 2024; Liu et al., 2024c), typically rank existing content based on user interaction history and contextual information, delivering the top-ranked content. However, constrained by the limited diversity of available content, they often fall short of meeting users’ diverse needs (Xu et al., 2024; Yu et al., 2019; Wang et al., 2023a), motivating the emergence of personalized content generation across various domains.
@@ 5. Conclusion and Future Work
Appendix A Domain Applications of Pigeon
Pigeon empowers LMMs with the capability to generate personalized images, which is applicable in various scenarios such as personalized stickers on social media platforms like Twitter and personalized movie posters on platforms like Netflix (see demonstration in Section 3). Beyond these, we showcase the potential of Pigeon in other representative domains.
Appendix B Datasets
For the sticker scenario, we exclude low-quality themes or those with fewer than six stickers, constructing user interaction sequences where each user interacts with a single theme. For the movie scenario, we adopt the small version of the dataset, retaining user interactions with ratings of four or higher, sorted by the timestamps. We apply a sliding window of six interactions, moving one step at a time to create data samples for each user in both scenarios. Each sample treats the first five interactions as the user history images and the last as the target image. We split the samples into training, validation, and testing sets with a ratio of 8:1:1. In the sticker testing set, we randomly select one sticker from a different theme than the user history as the reference image, while in the movie poster scenario, the target image is used as the reference. Dataset statistics are summarized in Table 4, where each “sample” consists of user-interacted history images and one reference image.
Table 4. Overview of dataset statistics.
@@ Appendix B Datasets
Appendix C Implementation Details of Pigeon
In Pigeon, the learning rate is set to $1e^{-5}$ and $5e^{-6}$ for the first and second stage alignment, respectively. The history mask ratio $\alpha_{h}$ is fixed at $0.2$. During inference, we select the optimal reference mask ratio $\alpha_{r}\in\{0.0,0.1,\dots,1.0\}$ for each reference image by averaging the history CIS and reference CS.
All experiments are conducted using a single NVIDIA-A100 GPU. As shown in Table 5, while the total number of parameters in Pigeon is substantial, the trainable components represent only a small fraction, leading to relatively low computational overhead. The training process costs about 20 hours and 5 hours for the first and second stage alignment, respectively. For inference, each sample takes about 7 seconds for LaVIT and 9 seconds for SDXL.
Table 5. Model parameters and trainable ratio of Pigeon.
Parameters Trainable Ratio
Total: LaVIT + SDXL 11,468,249,325 -
Trainable Mask Generator 100,726,784 0.878%
Adapter Layer 3,150,336 0.027%
LoRA 4,194,304 0.037%
Appendix D Human Evaluation
In the human evaluation, we conduct binary-choice tests across both sticker and movie scenarios, each consisting of 50 cases. To ensure diversity, the sticker cases involve 49 distinct themes and 6 different emotion labels, including anger, fear, happiness, neutral, sadness, and surprise. On the other hand, the movie cases include 21 different movie genres, such as action, animation, horror, romance, and sci-fi. As shown in Table 2, we perform a total of 10 binary tests, comparing Pigeon with Grd, TI, and PMG. For each test, 50 participants were recruited for evaluation. The qualitative results validate the effectiveness of Pigeon in personalized image generation across these diverse scenarios and contexts.
Appendix E Analysis of Preference Reward Strategy
To evaluate the impact of different preference reward strategies during the second-stage alignment, we conduct additional experiments to compare the strategy outlined in Eq.(8) with the approach proposed in (Shen et al., 2024). From the results shown in Table 6, we observe the following key findings:
  • The second-stage preference alignment significantly improves personalization, despite a slight decline in semantic alignment. This highlights the effectiveness of DPO in better aligning the generation process with user preferences.
  • The reward strategy employed by Pigeon achieves superior effectiveness in enhancing personalization, with only a minor and acceptable trade-off in semantic alignment. This indicates that the strategy successfully prioritizes user-specific preferences while maintaining a reasonable degree of semantic consistency.
Table 6. Effect of different preference reward strategies during the second-stage alignment.
#Sticker Personalization Semantic Alignment
[…]
#Movie Personalization Semantic Alignment
[…]
Appendix F Case Study